OpenStudy (anonymous):

What is the simplified form of the quantity 4 z squared minus 16z plus 15 over the quantity 2 z squared minus 11 z plus 15 ?

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

OpenStudy (anonymous):

(4z^2-16z+15)/(2z^2-11z+15) factor both the top and bottom and see if anything cancels out

OpenStudy (anonymous):

Those aren't factorable, are they? @CalebBeavers

OpenStudy (anonymous):

yep they are. =p

OpenStudy (anonymous):

what number goes into 4, 16, and 15 evenly?

OpenStudy (anonymous):

This is what ya do. 4z^2-16z+15 take coefficient of z^2 and multiply it by the last number so 4*15=60 now, what two numbers multiply together to give you 60 but add or subtract to give you -16?

OpenStudy (anonymous):

idk...

OpenStudy (anonymous):

Think about it. -10*-6=60 -10-6=-16 so 4z^2-6z-10z+15 factor first half then second half

OpenStudy (anonymous):

2x(2x-3)-2(2x-3)?

OpenStudy (anonymous):

2z(2z-3)-5(2z-3)

OpenStudy (anonymous):

so (2z-5)(2z-3) now factor the bottom

OpenStudy (anonymous):

(z-3)(2z-5)?